OpenStudy (anonymous):

8^x-4=64(3^x) can anyone help

OpenStudy (anonymous):

\[8^{x-4}\] or \[8^x-4\]?

OpenStudy (anonymous):

8^x-4 logarithmic function

OpenStudy (anonymous):

so the x - 4 is all in the exponent?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I get to the point x=log 64+4log 8/log 8-log3 how do I divide this?

OpenStudy (anonymous):

ok then that makes it easier. there may be some snap way to do this but i think you simply have to take the log of both sides to get \[(x-4)\ln(8)=\ln(64\times 3^x)=\ln(64)+x\ln(3)\] and then use algebra

OpenStudy (anonymous):

ok I type this in my calculator and get the wrong answer the x=12.721 I dont get this

OpenStudy (anonymous):

\[\ln(8) x -4\ln(8)=\ln(64)+\ln(3)x\] \[3\ln(2)x-12\ln(2)=6\ln(2)+\ln(3)x\] first then the algebra

OpenStudy (anonymous):

so you change the base

OpenStudy (anonymous):

\[3\ln(2)x-\ln(3)x=18\ln(2)\] \[x=\frac{18\ln(2)}{3\ln(2)-\ln(3)}\]

OpenStudy (anonymous):

ok how do you divide this?

OpenStudy (anonymous):

no i rewrote as follows: \[\ln(8)=\ln(2^3)=3\ln(2)\] and \[\ln(64)=\ln(2^6)=6\ln(2)\]

OpenStudy (anonymous):

ok thanks i think i get it:)

OpenStudy (anonymous):

i am not sure what you mean by "how do i divide" i don't know any of these numbers so if you want a decimal approximation i would put it in the calculator

OpenStudy (anonymous):

yw